OpenStudy (anonymous):

What is the slope of a line that is perpendicular to 6x + 3y = –9?

OpenStudy (anonymous):

Slope intercept form is y = mx + b, so you need to solve for y. Remember that whatever you do on one side of an equation, you must do on the other side. 6x - 3y = 9 Subtract 6x from each side -3y = -6x + 9 Divide each side by -3 -3y/-3 = -6x/-3 + 9/-3 y = 2x - 3

OpenStudy (mertsj):

What is the slope of the given line?